OpenStudy (anonymous):

What is the sum of all square numbers from 0 to 101?

OpenStudy (anonymous):

\[1^{2}+2^{2}+...+10^{2}=10\times(10+1)\times(2\times10+1)\div6=385\]

OpenStudy (anonymous):

ok thanks!

OpenStudy (experimentx):

\( 1^2+2^2+3^2+ ... + n^2 = \frac{n(n+1)(2n+1)}{6} \) put n=10